# rtboxstyle linecolor=black, backgroundcolor=lightlight-gray, leftmargin=0cm, rightmargin=0cm, skipabove=4pt, skipbelow=4pt, linewidth=0.4pt, innertopmargin=3pt, innerbottommargin=3pt, innerrightmargin=6pt, innerleftmargin=6pt, roundcorner=2

### v1

@@ 5.2 Fidelity Evaluation
where $P$ and $Q$ denote reference and generated graph distributions, $x,x^{\prime}$ are samples from $P$, and $y,y^{\prime}$ from $Q$.
In the experimental results shown in Table II, ProvSyn consistently outperformed the baselines across all datasets. In Nodlink dataset, ProvSyn achieved values an order of magnitude lower than the baselines on all metrics, indicating superior structural fidelity. In Cadets, Theia, and Trace datasets, ProvSyn also maintained a clear lead, particularly on label-related metrics. In contrast, baseline methods exhibit inconsistent performance patterns across the four datasets, suggesting fundamental limitations in structural generation capabilities regardless of the scale of model parameters. This finding underscores the necessity of an additional heterogeneous graph generation network, which produces structurally superior graphs compared to those generated solely by an LLM.
Texutual Evaluation. To evaluate the accuracy of generated text attributes, we use standard NLP metrics that capture different types of similarity. These include: (1) BLEU [51], which measures n-gram precision; (2) GLEU [49], which balances precision and recall; and (3) ROUGE [40], which focuses on recall by assessing how much of the reference text is covered.
For each node type, we construct a reference corpus $\mathcal{R}_{t}=\{r_{1},r_{2},...,r_{n}\}$ by aggregating all valid node names of type $t$ from the real-word provenance graph. When evaluating a generated name $g_{i}$, we compute its similarity with every reference $r_{j}\in\mathcal{R}_{t}$ through parallel metric calculations, then select the maximum similarity score as the final assessment:
$S(g_{i})=\max_{r_{j}\in\mathcal{R}_{t}}\text{sim}(g_{i},r_{j})$
The choice of maximum similarity over average similarity is driven by the following considerations. Focusing on the highest match ensures that the generated text is evaluated based on its optimal alignment with the reference corpus, thereby highlighting its capacity to produce accurate and relevant attributes. The complete procedure is presented in Algorithm 2 in Appendix B.
@@
In our experiments, we used the Python NLTK library to compute the BLEU, GLEU, and ROUGE scores for the process and file nodes. We didn’t evaluate the text quality of network nodes because their names are purely numerical, making natural language metrics inappropriate. Additionally, our analysis shows that the network node names generated by both the baselines and ProvSyn conform to the IPv4 address format.
The results in Table III show that the node names generated by the baselines generally achieve scores below 0.1, indicating that large language models lack prior knowledge of the names of operating system entities. In contrast, ProvSyn was exposed to real entity names from audit logs during the finetuning stage, enabling a smaller 3B model to generate more accurate and realistic names. Interestingly, we observed that Llama-8B performed relatively well in the Cadets datasets. Further analysis revealed that many of the node names generated by Llama were directly copied from the examples in the prompt, even though the corresponding log entries did not make sense.
@@
Temporal Evaluation. A provenance graph is a directed graph in which directed sequences also represent temporal sequences of events occurring in chronological order. For synthetic provenance graphs, accurately preserving the temporal structure of event sequences is essential for ensuring fidelity. To evaluate the temporal consistency of synthetic provenance graphs, we employ two metrics: Longest Common Subsequence (LCS) [7] and Dynamic Time Warping (DTW) [48].
@@
LCS measures the length of the longest sequence of events that appears in both the real and synthetic graphs in the same order. LCS is a strict matching metric—it requires exact alignment of events and their temporal order without allowing for any deviations or shifts. DTW, on the other hand, is a more flexible metric that allows for non-linear alignments between sequences. Unlike strict matching methods, DTW allows sequences to be stretched or compressed along the time axis, thereby accommodating local variations. This flexibility makes DTW particularly suitable for scenarios where the synthetic graph captures the general pattern of the real sequence but exhibits slight shifts or inconsistencies.
In our experiments, we used depth-first search to extract sequences from the graphs. Each sequence consisted of node types and edge types, e.g. (process, clone, process, read, file). For LCS computation, we used a dynamic programming approach, and for DTW, we used the dtw library in Python. To compare synthetic graphs with the real graphs in the test set, we adopted a strategy similar to Algorithm 2 in Appendix B: each sequence from the synthetic graph was compared with all sequences from the real graph, and the best match was recorded as its score. The overall graph score was then computed by aggregating the scores of all sequences in the synthetic graph. The results in Table IV show that ProvSyn achieves higher LCS scores and lower DTW values than the baselines across all four datasets, indicating its advantage in preserving temporal consistency.
TABLE III: Comparison of textual quality across different models, types and datasets. Lower values indicate better performance.
Dataset Type Model BLEU GLEU ROUGE-L-F
Nodlink Process GPT-4o 0.05 0.11 0.23
Claude-3.7 0.08 0.15 0.27
Doubao-pro 0.02 0.03 0.09
Qwen2.5-7B 0.07 0.15 0.29
Llama3.1-8B 0.12 0.17 0.33
ProvSyn 0.89 1.00 0.99
File GPT-4o 0.09 0.11 0.29
Claude-3.7 0.10 0.11 0.29
Doubao-pro 0.06 0.11 0.25
Qwen2.5-7B 0.11 0.17 0.41
Llama3.1-8B 0.26 0.26 0.37
ProvSyn 0.31 0.32 0.44
Cadets Process GPT-4o 0.24 0.24 0.24
Claude-3.7 0.17 0.17 0.17
Doubao-pro 0.03 0.03 0.03
Qwen2.5-7B 0.03 0.03 0.04
Llama3.1-8B 1.00 1.00 1.00
ProvSyn 1.00 1.00 1.00
File GPT-4o 0.01 0.01 0.20
Claude-3.7 0.03 0.03 0.14
Doubao-pro 0.01 0.01 0.14
Qwen2.5-7B 0.02 0.02 0.16
Llama3.1-8B 0.00 0.00 0.03
ProvSyn 0.31 0.32 0.60
Theia Process GPT-4o 0.08 0.08 0.09
Claude-3.7 0.04 0.04 0.05
Doubao-pro 0.01 0.01 0.03
Qwen2.5-7B 0.19 0.19 0.20
Llama3.1-8B 0.00 0.00 0.00
ProvSyn 0.76 0.89 0.47
File GPT-4o 0.03 0.03 0.21
Claude-3.7 0.05 0.05 0.16
Doubao-pro 0.01 0.01 0.11
Qwen2.5-7B 0.00 0.00 0.29
Llama3.1-8B 0.00 0.00 0.35
ProvSyn 0.37 0.44 0.48
Trace Process GPT-4o 0.02 0.02 0.02
Claude-3.7 0.02 0.02 0.02
Doubao-pro 0.00 0.00 0.00
Qwen2.5-7B 0.02 0.02 0.02
Llama3.1-8B 0.00 0.00 0.00
ProvSyn 0.50 0.50 0.50
File GPT-4o 0.00 0.00 0.03
Claude-3.7 0.00 0.00 0.03
Doubao-pro 0.00 0.00 0.01
Qwen2.5-7B 0.00 0.00 0.02
Llama3.1-8B 0.00 0.00 0.00
ProvSyn 0.00 0.00 0.67
@@
Graph Embedding Evaluation. To evaluate the fidelity of synthetic graphs from an embedding perspective, we employ graph representation learning techniques to project both synthetic and real graphs into a shared vector space, and measure their similarity using cosine similarity. Specifically, we adopt two representative embedding methods: DeepWalk [52] and Doc2Vec [35].
DeepWalk is a foundational algorithm in graph representation learning. It learns node embeddings by simulating truncated random walks on the graph and treating the resulting node sequences as sentences for a language modeling task. Given a graph $G$, we perform $N$ random walks of length $L$ starting from each node, generating a corpus $\mathcal{W}=\{w_{1},w_{2},\ldots,w_{N}\}$, where each $w_{i}$ is a walk sequence of node identifiers. These sequences are then used to train a Word2Vec model, producing node-level embeddings. The graph-level embedding $\mathbf{v}_{\text{deepwalk}}$ is computed as the mean of all node embeddings:
$\mathbf{v}_{\text{deepwalk}}=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\text{deepwalk}(v)$
Notably, DeepWalk does not consider textual attributes of nodes.
Unlike DeepWalk, Doc2Vec leverages node name information to capture the textual semantics of the graph. It treats each graph as a document, which is constructed from a collection of random walks. For a given graph $G$, we perform $N$ random walks of length $L$, resulting in a sequence set $\mathcal{W}=\{w_{1},w_{2},\ldots,w_{N}\}$, where each $w_{i}$ is a sequence of node names. These sequences are used to train a Doc2Vec model, yielding a graph-level embedding $\mathbf{v}_{\text{doc2vec}}$:
$\mathbf{v}_{\text{doc2vec}}=\text{doc2vec}(\mathcal{W})$
@@
A higher similarity score indicates that the synthetic graph exhibits greater fidelity to the real graph in the embedding space.
For DeepWalk-based similarity, we use the Word2Vec model from the Gensim library, and for Doc2Vec-based similarity, we use Gensim’s Doc2Vec model. In our experiments results shown in Table IV, ProvSyn outperforms the baselines on the Nodlink, Cadets, and Trace datasets, while performing slightly worse than GPT and Claude on Theia. The results show that ProvSyn maintains stable embedding similarity, with DeepWalk scores above 0.80 and Doc2Vec scores above 0.65 across four datasets. In comparison, the baselines exhibit greater variation across different datasets.
TABLE IV: Comparison of Temporal and Embedding metrics across different datasets and models. For LCS, higher is better; for DTW, lower is better; for DeepWalk and Doc2Vec, higher is better.
Dataset Model Temporal Metrics Embedding Metrics
LCS DTW DeepWalk Doc2Vec
Nodlink GPT-4o 2.44 1.54 0.66 0.69
Claude-3.7 3.03 1.33 0.83 0.40
Doubao-pro 2.64 0.89 0.60 0.59
Qwen2.5-7B 2.84 0.56 0.58 0.32
Llama3.1-8B 2.88 0.83 0.50 0.16
ProvSyn 3.30 0.41 0.89 0.74
Cadets GPT-4o 2.99 0.48 0.70 0.40
Claude-3.7 2.87 1.20 0.73 0.39
Doubao-pro 2.68 3.62 0.64 0.71
Qwen2.5-7B 2.13 24.30 0.79 0.25
Llama3.1-8B 2.96 0.48 0.48 0.67
ProvSyn 3.01 0.21 0.90 0.82
Theia GPT-4o 2.39 0.65 0.89 0.53
Claude-3.7 1.80 0.74 0.82 0.70
Doubao-pro 2.78 1.43 0.56 0.39
Qwen2.5-7B 1.36 8.90 0.74 0.48
Llama3.1-8B 1.04 8.94 0.39 0.66
ProvSyn 3.02 0.00 0.81 0.65
Trace GPT-4o 2.16 0.46 0.70 0.27
Claude-3.7 2.15 0.54 0.67 0.41
Doubao-pro 2.09 0.40 0.57 0.08
Qwen2.5-7B 2.10 0.47 0.75 0.09
Llama3.1-8B 2.08 4.45 0.71 -0.01
ProvSyn 2.80 0.10 0.81 0.75
5.3 Semantic Correctness Evaluation

### v2

@@ 4.3 Fidelity Evaluation Framework
Textual Evaluation. To evaluate the fidelity and diversity of generated text attributes, we adopt two standard NLP metrics: (1) GLEU [mutton2007gleu], which evaluates fidelity through n-gram overlap and is effective for short sequences; and (2) Distinct-1 [li2016diversity], which measures lexical diversity by computing the proportion of unique unigrams in the generated text.
To adapt GLEU for graph textual attributes, we employ a reference-based greedy matching approach [rus2012optimal, kusner2015word, zhang2019bertscore] that focuses on the highest match to avoid dilution from irrelevant references. For each node type $t$, we construct a reference corpus $\mathcal{R}_{t}=\{r_{1},r_{2},\ldots,r_{n}\}$ by aggregating all valid node names from the real-world provenance graph. For a set of generated names $\mathcal{G}_{t}=\{g_{1},g_{2},\ldots,g_{|\mathcal{G}_{t}|}\}$, we compute similarities between each $g_{i}$ and all $r_{j}\in\mathcal{R}_{t}$, select the maximum similarity via greedy matching, and average these values:
$S_{t}=\frac{1}{|\mathcal{G}_{t}|}\sum_{g_{i}\in\mathcal{G}_{t}}\max_{r_{j}\in\mathcal{R}_{t}}\text{GLEU}(g_{i},r_{j}).$
@@
We didn’t evaluate the text quality of network nodes because their names are purely numerical, making natural language metrics inappropriate. Additionally, our analysis shows that the network node names generated by both the baselines and ProvSyn conform to the IPv4 address format.
@@
Temporal Evaluation. A provenance graph is a directed graph in which directed sequences also represent temporal sequences of events occurring in chronological order. For synthetic provenance graphs, accurately preserving the temporal structure of event sequences is essential for ensuring fidelity. To evaluate the temporal consistency of synthetic provenance graphs, we employ two metrics: Longest Common Subsequence (LCS) [bergroth2000survey] and Dynamic Time Warping (DTW) [muller2007dynamic].
@@
LCS measures the longest event sequence appearing in both real and synthetic graphs in identical order. As a strict matching metric, LCS requires exact event alignment and temporal ordering without deviations. Conversely, DTW offers flexibility by allowing non-linear sequence alignments. Unlike strict methods, DTW enables sequences to be stretched or compressed along the time axis to accommodate local variations. This makes DTW suitable when synthetic graphs capture general patterns but exhibit slight temporal shifts.
@@
Embedding Evaluation. To evaluate the fidelity of synthetic graphs from embedding perspective, we employ graph representation learning techniques to project both synthetic and real graphs into a shared vector space, and measure their cosine similarity. Specifically, we adopt two representative embedding methods: NetLSD [tsitsulin2018netlsd] and Graph2Vec [narayanan2017graph2vec].
NetLSD creates graph-level embeddings by summarizing a graph’s multi-scale structural properties using the heat kernel trace across different time scales. Its embedding vector is: $\mathbf{v}_{\text{netlsd}}=[\mathrm{tr}(e^{-t_{1}\mathbf{L}}),\mathrm{tr}(e^{-t_{2}\mathbf{L}}),\ldots,\mathrm{tr}(e^{-t_{K}\mathbf{L}})]$, where $\mathbf{L}$ is the graph’s Laplacian matrix, $t_{k}$ are diffusion time scales, and $\mathrm{tr}(\cdot)$ computes the matrix trace.
Unlike NetLSD, Graph2Vec leverages node text attributes to model semantics and learns embeddings by modeling each graph as a corpus of WL-relabeled random walk sequences and applying document embedding techniques. Its embedding is defined as: $\mathbf{v}_{graph2vec}=\text{doc2vec}(\mathcal{W})$, where $\mathcal{W}=\{w_{1},w_{2},\ldots,w_{N}\}$ represents the set of random walk sequences generated from the WL-refined graph, and the doc2vec model produces the final graph embedding.
@@
A higher similarity indicates that the synthetic graph exhibits greater fidelity to the real graph in the embedding space.
Semantic Evaluation. Evaluating semantic correctness is essential for assessing the fidelity of synthetic provenance graphs, as it indicates whether the generated graphs accurately reflect the underlying logic and intent of real-world system behaviors. This evaluation examines whether the relationships among entities and the sequences of interactions conform to meaningful and valid operational patterns. Despite its importance, evaluating semantic correctness remains difficult due to the lack of standardized metrics and the inherent complexity in defining ground-truth semantics for diverse and dynamic system activities.
To evaluate semantic correctness, we formally define a semantic unit in the graph as a triplet: the starting node name $s$, edge type $e$, and ending node name $t$. This unit is represented as $(s,e,t)$, where $s,t\in\mathcal{N}$ (the set of all node names) and $e\in\mathcal{E}$ (the set of all edge types). Concatenating these elements forms a natural language sentence, denoted $\text{Sentence}(s,e,t)$. This sentence captures the semantics of the interaction between entities, expressing that the source node $s$ performs an action of type $e$ on the target node $t$.
To assess the semantic correctness of $\text{Sentence}(s,e,t)$, we employ a BERT(Bidirectional Encoder Representations from Transformers) [devlin2019bertpretrainingdeepbidirectional]-based encoder to obtain its embedding. This embedding captures the semantic representation of the interaction described by the triplet and serves as the basis for subsequent evaluation. Formally, let $\text{BERT}:\mathcal{S}\to\mathbb{R}^{d}$ be the encoding function that maps a sentence $\text{Sentence}(s,e,t)\in\mathcal{S}$ (the set of all possible semantic triplets in sentence form) to a dense vector $\mathbf{v}\in\mathbb{R}^{d}$. The embedding is computed as $\mathbf{v}=\text{BERT}(\text{Sentence}(s,e,t))$.
However, embedding of $\text{Sentence}(s,e,t)$ captures only pairwise semantic relationships, resulting in a localized evaluation. To achieve graph-level assessment, we employ a multi-layer Graph Attention Network (GAT) [velickovic2017graph]. This architecture aggregates neighborhood information through stacked self-attention layers, where each layer updates node representations by adaptively weighting features from adjacent nodes. We employ BERT to represent the text attributes of individual nodes as the initial node embeddings. Formally, for each node $v$ with node name $T_{v}$, its initial embedding $\mathbf{h}_{v}^{(0)}$ is derived via BERT: $\mathbf{h}_{v}^{(0)}=\text{BERT}(T_{v})$. Let $\text{GAT}:\mathcal{V}\to\mathbb{R}^{d^{\prime}}$ denote the encoder that maps a node to its global representation in $\mathbb{R}^{d^{\prime}}$ using these initial embeddings. For a semantic unit $(s,e,t)$, we obtain the global embeddings of both the starting node $s$ and the ending node $t$:
We then concatenate the global embeddings $\mathbf{u}_{s}$ and $\mathbf{u}_{t}$ with the local embedding $\mathbf{v}$ of the semantic unit to form a comprehensive representation. This comprehensive representation $\mathbf{z}=[\mathbf{v};\mathbf{u}_{s};\mathbf{u}_{t}]$ captures both the local and global context of the semantic unit.
To quantify semantic correctness, we construct a reference semantic corpus from the real-world provenance graph by extracting all valid semantic units $(s,e,t)$ and encoding them using the same procedure described above, yielding a set of reference representations $\mathcal{Z}_{\text{ref}}=\{\mathbf{z}_{1},\mathbf{z}_{2},\ldots,\mathbf{z}_{m}\}$.
Given a set of generated semantic units $\mathcal{Z}_{\text{gen}}$, we evaluate semantic correctness using the same greedy matching strategy [rus2012optimal, kusner2015word, zhang2019bertscore] as in the textual evaluation. Specifically, for each generated representation $\mathbf{z}\in\mathcal{Z}_{\text{gen}}$, we compute its maximum similarity with the reference corpus and then average over all generated units:
$S_{\text{sem}}=\frac{1}{|\mathcal{Z}_{\text{gen}}|}\sum_{\mathbf{z}\in\mathcal{Z}_{\text{gen}}}\max_{\mathbf{z}_{j}\in\mathcal{Z}_{\text{ref}}}\text{sim}(\mathbf{z},\mathbf{z}_{j}).$
5 Evaluation
To thoroughly evaluate ProvSyn, we address the following research questions. Experiments were performed on a machine with Intel Xeon Platinum CPU, 512 GB RAM, NVIDIA RTX A6000 GPUs, and Ubuntu 22.04.3 LTS. The research questions are:
RQ1: What is the fidelity of synthetic provenance graphs in terms of structural, textual, temporal, embedding, and semantic characteristics?
RQ2: Can synthetic provenance graphs mitigate class imbalance in the dataset?
RQ3: When synthetic provenance graphs are used for training set augmentation in downstream intrusion detection tasks, what is the performance of the detection task?
RQ4: What are the time and resource costs of generating synthetic provenance graphs?
Further aspects of our study, including hyperparameter analysis, ablation study are discussed in the Appendix D and  E.
5.1 Experimental Setup
Datasets. We evaluate the effectiveness of ProvSyn on three public datasets: the DARPA Engagement 3 dataset [darpa-E3], the DARPA Engagement 5 dataset [darpa-E5], and the OpTC dataset [optc-data], utilizing the labels from Orthrus[jiang2025orthrus]. From each of these datasets, we select two subsets, resulting in a total of six datasets: Cadets-E3, Theia-E3, Clearscope-E5, Theia-E5, OpTC-H501, and OpTC-H201. The provenance graphs in these datasets contain three types of nodes and ten types of events, with the number of nodes ranging from 50k to 2,000k and the number of edges ranging from 100k to 8,000k. For each dataset, we use provenance data from the first day for training and fidelity evaluation, with the two sets being derived from distinct time windows and strictly separated.
Baselines. In our approach, we first use a heterogeneous graph generation model to construct the structure of the provenance graph and then employ an LLM to generate the textual attributes of the graph. In contrast, an alternative method directly calls an LLM to synthesize provenance logs and then converts these logs into a graph. This method does not require an additional heterogeneous graph generation model. We systematically evaluate and compare the provenance data synthesized by LLMs with the data synthesized by ProvSyn.
To ensure a competitive comparison, we selected a set of strong baseline models, including GPT-5.1 [openai-gpt51], Claude-Sonnet-4.5 [claude], DeepSeek-v3.2 [liu2025deepseek], and Qwen3-Max [yang2025qwen3]. These models are widely regarded as state-of-the-art LLMs at the time of evaluation. To generate provenance logs using LLMs listed above, we designed structured prompts that included task descriptions and illustrative examples. Each dataset was paired with a customized prompt, primarily differing in the examples provided. These examples contained five elements: source node type, source node name, destination node type, destination node name, and edge type. They were selected from real audit logs to ensure both representativeness and coverage of various node and edge types. For each dataset, we collected 100 responses from each LLM. After obtaining the responses, we filter out entries that contain node or edge types outside of the predefined type sets. The remaining entries were then used to construct graphs for comparison with graphs generated by ProvSyn. The prompt used in our experiments is provided in Appendix C (Figure 7).
Table 1: Comprehensive Comparison of MMD, Temporal, and Embedding Metrics. For MMD and DTW, lower is better. For LCS, NetLSD, and Graph2Vec, higher is better. Bold values indicate the best performance in each category.
Dataset Model MMD Metrics ($\downarrow$) Temporal Metrics Embedding Metrics ($\uparrow$)
Degree Orbit Edge L. Node L.+Deg LCS ($\uparrow$) DTW ($\downarrow$) NetLSD Graph2Vec
Cadets_E3 GPT-5.1 0.59 0.87 0.41 1.16 3.02 1.58 0.48 0.29
Claude-sonnet-4.5 1.01 0.96 0.35 1.28 4.07 1.62 0.16 0.33
DeepSeek-v3.2 1.47 0.88 0.72 1.45 3.03 1.69 0.03 0.34
Qwen3-max 1.36 0.94 0.10 1.28 3.00 0.00 0.03 0.29
ProvSyn 0.37 0.79 0.01 0.61 4.71 0.74 0.49 0.35
Theia_E3 GPT-5.1 0.28 0.37 0.17 0.60 3.14 0.53 0.34 0.36
Claude-sonnet-4.5 0.42 0.33 0.19 0.73 3.00 0.00 0.83 0.34
DeepSeek-v3.2 0.57 0.38 0.29 0.65 2.93 1.92 0.42 0.38
Qwen3-max 0.81 0.90 0.22 1.07 3.00 0.00 0.85 0.31
ProvSyn 0.07 0.19 0.07 0.29 3.19 0.94 0.53 0.47
ClearScope_E5 GPT-5.1 0.73 0.47 0.01 1.16 3.00 1.64 0.54 0.18
Claude-sonnet-4.5 0.61 1.01 0.02 1.11 3.00 1.71 0.60 0.19
DeepSeek-v3.2 0.56 0.49 0.01 0.93 3.00 1.64 0.41 0.19
Qwen3-max 0.57 0.59 0.00 0.64 3.00 1.66 0.62 0.19
ProvSyn 0.22 0.38 0.00 0.23 3.23 1.48 0.73 0.34
Theia_E5 GPT-5.1 0.66 0.60 0.12 0.67 3.03 0.45 0.42 0.37
Claude-sonnet-4.5 0.60 0.72 0.08 0.81 3.21 0.46 0.36 0.37
DeepSeek-v3.2 0.70 1.10 0.43 0.54 3.02 0.53 0.10 0.40
Qwen3-max 0.66 1.06 0.18 0.69 3.03 0.42 0.23 0.37
ProvSyn 0.13 0.29 0.03 0.20 3.02 0.35 0.53 0.35
OpTC_H201 GPT-5.1 0.65 0.75 0.51 0.62 4.09 2.98 0.28 0.23
Claude-sonnet-4.5 0.98 1.22 1.93 1.77 3.50 3.04 0.04 0.21
DeepSeek-v3.2 0.59 0.96 1.17 1.15 2.86 3.32 0.01 0.28
Qwen3-max 0.55 0.79 1.00 1.05 2.76 3.36 0.01 0.29
ProvSyn 0.10 0.51 0.11 0.17 4.74 1.73 0.64 0.33
OpTC_H501 GPT-5.1 0.85 1.08 1.13 1.04 3.86 2.56 0.14 0.31
Claude-sonnet-4.5 0.45 1.09 0.92 1.17 3.90 2.74 0.03 0.26
DeepSeek-v3.2 0.77 1.19 0.93 1.52 2.45 3.71 0.01 0.37
Qwen3-max 0.67 1.09 1.54 1.38 2.64 3.30 0.01 0.37
ProvSyn 0.08 0.52 0.02 0.11 4.65 1.78 0.61 0.40
Metrics. For fidelity evaluation, we adopt the five-dimensional metrics introduced in Section 4.3. For class imbalance mitigation, we use three metrics: novelty, uniqueness, and normalized entropy. Novelty verifies that each generated graph is neither a subgraph of any training reference graph nor contains any reference graph as a subgraph, indicating the presence of novel structures. Uniqueness measures the proportion of generated graphs that are structurally distinct from one another, reflecting diversity. Normalized entropy measures label uncertainty and rescales it by the maximum entropy under the same number of classes, yielding values in the range $[0,1]$. Higher values indicate a more balanced label distribution. For IDS application, beyond the standard F-score, we report three robust metrics for comprehensive evaluation: Balanced Accuracy (B.Acc), which reduces the impact of imbalanced test data; the Matthews Correlation Coefficient (MCC) [jiang2025orthrus], which measures the correlation between predicted and true labels; and the Diagnostic Odds Ratio (DOR), which reflects diagnostic performance as the odds ratio of positive results for attacks versus normal behaviors.
Configurations. For training the graph generation model, we sampled 20,000 subgraphs from the real provenance graphs using RWR subgraph sampling strategy. Among them, 16,000 subgraphs were used for training, 2,000 for validation, and 2000 for test. The model was trained for 3,000 epochs with a batch size of 32 and a learning rate of 0.003. For the text attribute synthesizing task, we used Llama3.2-3B-Instruct [grattafiori2024llama] as the base model. We constructed a training set of 20,000 samples, consisting of 10,000 fully masked QA pairs and 10,000 partially masked QA pairs. During training, we applied the LoRA fine-tuning method, freezing most of the model parameters and updating only 0.81% of them. The model was trained for 60 steps with a learning rate of 2e-4. During inference, the temperature was set to 1.5, and the maximum number of generated tokens was set to 2048.
5.2 Fidelity Evaluation
Table 2: Comparison of textual quality across different models, node types, and datasets. Left: GLEU. Right: Distinct-1. The maximum value in each group is bolded.
Dataset Type GLEU Distinct-1
GPT Claude DeepSeek Qwen3 ProvSyn GPT Claude DeepSeek Qwen3 ProvSyn
Cadets-E3 Process 0.36 0.16 0.04 0.07 0.56 1.00 1.00 1.00 1.00 1.00
File 0.10 0.07 0.02 0.03 0.22 1.00 1.00 1.00 1.00 1.00
Theia-E3 Process 0.03 0.14 0.00 0.50 0.69 0.66 1.00 0.51 1.00 0.70
File 0.01 0.00 0.00 0.00 0.06 1.00 1.00 1.00 1.00 1.00
ClearScope-E5 Process 0.47 0.78 0.37 0.33 0.52 0.55 0.67 0.51 0.57 0.68
File 0.08 0.06 0.06 0.08 0.21 0.75 0.81 0.73 0.73 0.84
Theia-E5 Process 0.04 0.05 0.03 0.05 0.18 0.54 0.60 0.38 0.57 0.83
File 0.03 0.03 0.03 0.03 0.03 1.00 1.00 1.00 1.00 1.00
OpTC-H201 Process 0.21 0.33 0.19 0.19 0.55 0.31 0.38 0.22 0.40 0.46
File 0.17 0.15 0.15 0.19 0.30 0.70 0.94 0.95 0.99 0.71
OpTC-H501 Process 0.33 0.36 0.23 0.31 0.38 0.20 0.25 0.34 0.31 0.39
File 0.00 0.00 0.00 0.00 0.00 0.69 0.76 0.80 0.79 0.84
Structural Evaluation. Table 1 reports the structural fidelity between generated graphs and real data as measured by MMD metrics. Across all datasets, ProvSyn consistently achieves the lowest scores on degree, orbit, edge label, and node label &degree distributions, indicating closer alignment with the underlying graph structures. Notably, on the OpTC-H201 and OpTC-H501 datasets, ProvSyn reduces MMD scores by more than one order of magnitude compared to all baseline methods. In contrast, baseline models exhibit higher and less stable scores across datasets, suggesting limited ability to preserve structural fidelity. These results highlight the importance of incorporating a dedicated heterogeneous graph generation network to achieve faithful structural synthesis beyond standalone LLM capabilities.
Textual Evaluation. In our experiments, we used the Python NLTK library to compute the GLEU, and Distinct-1 scores for the process and file nodes. The results in Table 2 show that the GLEU score of node names generated by the baselines generally achieve scores below 0.20, indicating that large language models lack prior knowledge of the names of operating system entities. In contrast, ProvSyn was exposed to real entity names during the finetuning stage, enabling a smaller model to generate more accurate and realistic names. Regarding the Distinct-1 diversity metric, ProvSyn outperforms baseline models across most datasets and node types. In particular, for process nodes in ClearScope-E5 and Theia-E5, baseline models exhibit lower Distinct-1 scores, reflecting repeated or monotonous name generation. By contrast, ProvSyn produces more distinctive and diverse entity names, better reflecting the complex naming distributions of operating systems.
Temporal Evaluation. In our experiments, we used depth-first search to extract sequences from the graphs. Each sequence consisted of node types and edge types, e.g. (process, clone, process, read, file). For LCS computation, we used a dynamic programming approach, and for DTW, we used the dtw library in Python. Table 1 shows that across six datasets, ProvSyn achieves higher LCS scores and lower DTW values than the baselines in five datasets. In particular, for Cadets-E3, OpTC-H201, and OpTC-H501, ProvSyn attains LCS scores above 4.5, highlighting its effectiveness in maintaining temporal fidelity.
Embedding Evaluation. For graph embedding, we use the netlsd library for NetLSD and Gensim’s Doc2Vec for Graph2Vec. It is important to note that absolute similarity values are generally unreliable due to variations in graph size, embedding scale, and initialization; instead, relative comparisons within the same dataset provide meaningful interpretation. As shown in Table 1, across six datasets, ProvSyn consistently achieves the highest relative similarity scores in five datasets for both NetLSD and Graph2Vec. Notably, on the OpTC datasets, where all baselines have NetLSD scores below 0.30, ProvSyn maintains scores above 0.60, demonstrating its robustness in preserving embedding fidelity.
Semantic Evaluation In the experiment, prior to concatenating the BERT-derived node name embeddings with the GAT output embeddings, we employ a linear layer to project the BERT embeddings into the same dimensional space as the GAT outputs. This ensures prevents the concatenation process from being biased toward higher-dimensional components. Figure 3 shows that the provenance graphs generated by ProvSyn consistently outperform the baselines in semantic accuracy across all six datasets, demonstrating robustness to variations in data distribution. Notably, ProvSyn exhibits a substantial lead in the Theia-E3, ClearScope-E5, and OpTC-H501 datasets. For instance, on ClearScope-E5, ProvSyn achieves an accuracy of 0.98, significantly surpassing the best-performing baseline, Qwen3-max, which scores 0.74.